OpenStudy (anonymous):

Four more than twice a number is seven more than the number. What is the number?

OpenStudy (paki):

let the number be "x".... then 2x+4=x+7.... solve for "x"....

OpenStudy (anonymous):

add 4x to both sides 2x + 4x - 5 = -4x + 4x + 7 simplify 6x - 5 = 7 add 5 to both sides 6x - 5 + 5 = 7 + 5 simplify 6x = 12 divide both sides ...

OpenStudy (paki):

??????

OpenStudy (anonymous):

im lost

OpenStudy (paki):

no, its easy one... just combine the like terms...and solve...

OpenStudy (anonymous):

please help last one to answer

OpenStudy (paki):

post here...

OpenStudy (anonymous):

i cant paste it

OpenStudy (paki):

then,....?